# Vortices and Factorization

### v2

@@ Abstract
We review applications of factorization methods to the problem of finding stationary point vortex patterns in two-dimensional fluid mechanics. Then we present a new class of patterns related to periodic analogs of Schrodinger operators from the “even” bi-spectral family. We also show that patterns related to rational and soliton solutions of the KdV hierarchy constitute complete solution of the problem for certain classes of vortex systems.
@@ 1 Introduction, Stationary Vortex Configurations
In fluid mechanics, factorization methods found their applications in the theory of two-dimensional flows in two opposite limits: the infinite-viscosity limit of the free-boundary flows (see e.g. [26], [27] and references therein) and in the zero-viscosity limit. Here we review the latter, namely applications to the problem of finding classes of two-dimensional stationary vortex configurations in inviscid 2d flows.
Systems of point vortices are weak solutions of the two-dimensional Euler equation. Work on their classification started more than a century and a half ago and has been developed substantially over the past several decades (for general overview see e.g. [3, 4, 14, 31, 32] and references therein).
Motion of two-dimensional inviscid and incompressible liquid is described by the system of PDEs
$\frac{\partial\mathbf{v}}{\partial t}+(\mathbf{v}\nabla)\mathbf{v}=-\frac{\nabla P}{\rho},\quad\rho={\rm const},$ (1)
$\nabla\mathbf{v}=0,$ (2)
where ${\mathbf{v}}=\mathbf{v}(x,y,t)$, $\mathbf{v}=(u,w)$ is the two-dimensional velocity of the flow, and $P=P(x,y,t)$ stands for the pressure. Equation (1) is the Euler’s equation, while (2) is the continuity equation. In two dimensions, the vorticity, which is a curl of velocity, has a single component $\omega$:
@@
Taking curl of (1) we get
$\frac{d\omega}{dt}=\frac{\partial\omega}{\partial t}+(\mathbf{v}\nabla)\omega=0.$
@@
the strength (i.e. circulation) of each vortex ${\cal Q}_{i}$ is time independent and speed of $i$th vortex $(dx_{i}/dt,dy_{i}/dt)$ equals velocity of the flow averaged over a small circular domain centered at $(x_{i},y_{i})$. Since vorticity is a curl of velocity, for the system of point vortices we have
@@
where we used complex notations $z=x+{\rm i}y$. In (3) the full velocity field $\mathbf{v}=(u,w)$ is a sum of irrotational (i.e. zero-curl) background flow $\mathbf{V}=(U,W)$ and the flow created by superposition of $N$ point vortices. By (2, 3) $\mathbf{V}$ is a zero-divergence field. Since it is irrotational, it is also a gradient of a potential and as a consequence the potential is a harmonic function. In sections 1-8 we will consider unrestricted flows in the whole plane with velocities bounded at infinity. The only harmonic functions satisfying above conditions are linear functions and the background flow has constant velocity (uniform background flow). Using (3), we find the average velocity of the flow in a small disc centered at $z=z_{i}$, thus, obtaining the following equations of motion of vortices in the uniform background flow
@@
where $k$ is a complex constant.
Classification of the stationary patterns of vortices is then reduced to the study of solution of system of $N$ algebraic equations
@@
This system also has the 2d electrostatic interpretation: (5) is nothing, but an equilibrium condition of $N$ point electric charges ${\cal Q}_{1},\dots,{\cal Q}_{N}$ interacting pair-wise through 2d Coulomb (logarithmic) potential and placed in the homogeneous electric field $\bar{k}$. The equilibrium configuration is a stationary point of the electrostatic energy
@@
The stationary patterns with $k\not=0$ are called translating equilibria (or translating configurations/patterns) because in the reference frame where the fluid rests at infinity, the whole pattern translates with constant velocity $(\text{Im}\,k,\text{Re}\,k)/(2\pi)$. Below we mainly deal with equilibria without background flow, i.e. $k=0$ equilibrium configurations (“static configurations/patterns”):
@@
This system of equations is invariant under rigid motions and scaling of the plane. From invariance of stationary value of electrostatic energy under scaling of the plane it follows that the system has solutions only if $\sum_{i<j}Q_{i}Q_{j}=0$. O’Neil showed [31] that for almost all values of $Q_{i}$, satisfying the above condition, there are exactly $(N-2)!$ distinct configurations, i.e. modulo rigid motions and scaling of the plane, distinct solutions of (6).
In the case $k\not=0$ necessary and sufficient condition of existence of solutions is $\sum_{i}{\cal Q}_{i}=0$. For almost all values of ${\cal Q}_{i}$ satisfying this condition number of distinct, modulo translations of the plane, solutions of (5) is $(N-1)!$.
There are, however, “resonant” values of ${\cal Q}_{i}$, where number of configurations is infinite. This paper mainly deals with such resonant cases. Sections 10 and 11 are devoted to periodic (“vortex street”) versions of such systems.
We also briefly consider non-resonant cases in relationship with factorization method. These are equilibrium patterns in non-uniform background flows from section 9.
@@ 2 Vortices, Polynomials and Rational Primitives
Let us consider static equilibria of $N=l+m$ charges (or vortices) of two species. Without lost of generality we can set values of first $l$ charges to -1 and those of remaining $m$ charges to $\Lambda$:
@@
Introducing polynomials in $z$
@@
we first consider a situation where all charges have distinct positions, so $p$ and $q$ do not have multiple or common roots. One can check that the static equilibrium condition (6) for this system of two species is equivalent to the following bi-linear differential equation for the above polynomials
@@
where prime denotes differentiation wrt $z$. The equivalence can be established by the partial fraction decomposition of the ratio of (8) and $pq$.
Note, that (8) holds not only when $p$ and $q$ do not have common or multiple roots, but also in special cases when such roots exist. It trivially follows from the partial fraction decomposition of the ratio of (8) and $pq$ that if a common root exists, then it cannot be the simple one in both $p$ and $q$: Multiplicities $M$ and $M^{\prime}$ of this root, in $p$ and $q$ respectively, are constrained by the following relation:
$(M-\Lambda M^{\prime})^{2}=M+\Lambda^{2}M^{\prime}.$
In situation when a common root emerges as a limit of a generic solution, when $M$ vortices of strength $-1$ and $M^{\prime}$ vortices of strength $\Lambda$ collapse into a single vortex of strength $\Lambda M^{\prime}-M$, the limiting vortex patterns are called the collapse configurations. There are also exist common root solutions that are not collapse solutions. For example, Demina and Kudryashov [16] presented a family of such solutions using the fact that if a pair $p$, $q$ solves (8) for some $\Lambda$, then another pair $pq^{\Lambda(\Lambda-1)/2}$, $q^{\Lambda(\Lambda+1)/2}$, that corresponds to the same vortex pattern as the first one, satisfies
$p^{\prime\prime}q-2p^{\prime}q^{\prime}+pq^{\prime\prime}=0.$ (9)
@@
The above equation, which is the $\Lambda=1$ specification of (8), is called the Tkachenko equation [38] in literature on the vortex patterns [^1] Eq. (9) first appeared in work of Burchnall and Chaundy [7] in different context.. Its generic (i.e. no-common root) solutions describe equilibria of two species of vortices (charges) of equal magnitude but opposite sign. Bartman [9] identified the Tkachenko equation with the recurrence relation (16) for the Adler-Moser polynomials. As follows from works by Burchnall and Chaundy [7] (see the next Section) any generic polynomial solution of the Tkachenko equation is a pair of the Adler-Moser polynomials [1]
$p(z)=P_{n-1}(z),\quad q(z)=P_{n}(z).$ (10)
@@
In this way we get complete classification of static configuration of system of vortices (charges) of strengths ±$\pm 1$. The degree of the $n$th Adler-Moser polynomial equals $n(n+1)/2$. Therefore, for this system equilibrium is only possible if the number of negative and positive charges equals $l=n(n-1)/2$ and $m=n(n+1)/2$ respectively. The equilibrium coordinates of charges $z_{i}$ of each specie are roots of consecutive Adler-Moser polynomials. The $n$th Adler-Moser polynomial $P_{n}$ is also a function of $n$ free parameters, and the space of all possible equilibria of $N=l+m=n^{2}$ charges is a $n$-dimensional complex subspace of the $n^{2}$ dimensional complex space.
@@
Details on Adler-Moser polynomials will be given in the following sections. Here, we note that these polynomials were first found by Burchnall and Chaundy [7], later re-discovered in context of the Hadamard problem [23] by Lagnese and Stellmacher, and then by Adler and Moser as polynomial $\tau$-functions corresponding to rational solutions of the KdV hierarchy of integrable PDEs [1]. Burchnall and Chaundy [7] constructed these polynomials as solutions of the Tkachenko equation (9) by considering it as a linear second order ODE for $p$, where coefficients of ODE are determined by some polynomial $P_{n}=q$. Then the requirement that its two linearly independent solutions $p=P_{n-1}$ and $p=P_{n+1}$ are both polynomial is equivalent to rationality in $z$ of two indefinite integrals (rational primitives)
@@
It turned out that, apart from the Adler-Moser case there is another case with $\Lambda=2$ when there exists an infinite chain of polynomial solutions (8) that can be constructed with help of rational primitives [9, 24].
@@ 3 $\Lambda=1$ and $\Lambda=2$ Chains: Main Sequences
In order to build a recursive chain we require that both $p_{n-1}$ and $p_{n}$ be polynomial in $z$, so the primitive
@@
must be rational. Alternatively, considering (8) as an ODE for $q$, we come to the requirement of rationality of
@@
It is clear that pair of the above rationality conditions requires that both $2\Lambda$ and $2/\Lambda$ be integer, which is possible only if
@@
The case $\Lambda=1$ has already been mentioned and corresponds to the Adler-Moser polynomials. Since equation (8) is invariant under the involution $\Lambda\leftrightarrow 1/\Lambda$, $p\leftrightarrow q$, the cases $\Lambda=1/2$ and $\Lambda=2$ are equivalent, and, without loss of generality, both can be represented by the $\Lambda=2$ case.
Now, let us take $p$ and $q$ that do not have common/multiple roots and factorize $p(z)$ as $p(z)=(z-z_{i})\tilde{p}(z)$, where $z_{i},i\leq l$ is a root of $p$ (see definition (7)). Rationality of primitive (11) implies an absence of a simple pole at $z=z_{i}$, which is equivalent to the condition
$\frac{\partial}{\partial z}\left(\frac{q(z)^{2\Lambda}}{\tilde{p}(z)^{2}}\right)_{z=z_{i}}=0.$
Since $\tilde{p}(z_{i})=p^{\prime}(z_{i})$ and $\tilde{p}^{\prime}(z_{i})=p^{\prime\prime}(z_{i})/2$, we get the following equation
$p^{\prime\prime}(z_{i})q(z_{i})-2\Lambda p^{\prime}(z_{i})q^{\prime}(z_{i})=0$
which is clearly satisfied by solutions of (8). Similarly, rationality of (12) also follows from (8). It is worth mentioning that the converse statement also holds, i.e. the bilinear equation (8) follows from rationality of (11) and (12) (for details see [24]).
Thus, once a pair of polynomials $p,q$ without multiple and common roots satisfying (8,13) is found, one can build an infinite chain of polynomial solutions of (8) with the help of recursive procedure for $p_{n}$ and $q_{n}$ mentioned at the beginning of this section.
In more detail: Since $p=z^{l}+\dots$ and $q=z^{m}+\dots$ from the leading term of (8) we get the diophantine equation connecting $l$ and $m$
@@
Its solutions are
$l=l_{i},\quad m=m_{i},\qquad{\rm or}\qquad l=l_{i},\quad m=m_{i+1},\quad i\in\mathbb{Z},$ (14)
@@
Now, consider bilinear relation (8) with $p=p_{i}$ and $q=q_{i}$ as a linear second-order ODE with solution $q_{i}$. Its second linearly independent solution $q=q_{i+1}$ equals
@@
Similarly, considering $p_{i-1}$ and $p_{i}$ as linearly independent solutions, we get
@@
By freedom in choosing linearly independent solutions of (8) one can also write analogs of (18) and (19) for decreasing indices
@@
and one can generate $p_{i}$, and $q_{i}$ iteratively in either directions starting at some $i$.
Rewriting (18,19) or (20) in the differential form, we get the first-order differential recurrence relations
@@
In the case $\Lambda=1$, the above recurrence relations are the three-term relations for the Adler-Moser polynomials $P_{n}$ with $q_{i}=P_{2i}$ and $p_{i}=P_{2i+1}$:
@@
Examples of several first Adler-Moser polynomials are
$P_{0}=1,\quad P_{1}=z,\quad P_{2}=z^{3}+t_{1},\quad P_{3}=z^{6}+5t_{1}z^{3}+t_{2}z-5t_{1}^{2},\quad\dots,$ (23)
where $t_{i}$ are free parameters related to the KdV “times” [1] (we set parameter related to shift of $z$ to zero).
In the $\Lambda=2$ case we get

### v3

@@ Abstract
We review applications of factorization methods to the problem of finding stationary point vortex patterns in two-dimensional fluid mechanics. Then we present a new class of patterns related to periodic analogs of Schrodinger operators from the “even” bi-spectral family. We also show that patterns related to soliton solutions of the KdV hierarchy constitute complete solution of the problem for certain classes of vortex systems.
@@ 1 Introduction, Stationary Vortex Configurations
In fluid mechanics, factorization methods have found applications in the theory of two-dimensional flows in two opposite limits: the infinite-viscosity limit of the free-boundary flows (see, e.g., [33], [34]), and the zero-viscosity limit. In this review we focus on the latter, namely applications to the problem of finding classes of two-dimensional stationary vortex configurations in inviscid two-dimensional flows.
Systems of point vortices are weak solutions of the two-dimensional Euler equation. Work on their classification began more than a century and a half ago and has been substantially developed over the past several decades (for a general overview see, e.g., [4, 5, 6, 19, 39, 40]).
Motion of two-dimensional inviscid and incompressible liquid is described by the following system of PDEs
$\frac{\partial\mathbf{v}}{\partial t}+(\mathbf{v}\cdot\nabla)\mathbf{v}=-\frac{\nabla P}{\varrho},\quad\varrho={\rm const},$ (1)
$\nabla\cdot\mathbf{v}=0,$ (2)
where ${\mathbf{v}}=\mathbf{v}(x,y,t)$, $\mathbf{v}=(u,w)$ is the two-dimensional velocity of the flow, and $P=P(x,y,t)$ stands for the pressure. Equation (1) is the Euler’s equation, while (2) is the continuity equation. In two dimensions, the vorticity, which is the curl of the velocity field, has a single component $\omega$:
@@
Taking the curl of (1) we get
$\frac{d\omega}{dt}=\frac{\partial\omega}{\partial t}+(\mathbf{v}\cdot\nabla)\omega=0.$
@@
the strength (i.e., circulation) of each vortex ${\cal Q}_{i}$ is time independent, and the speed of the $i$th vortex $(dx_{i}/dt,dy_{i}/dt)$ equals the velocity of the flow averaged over a small circular domain centered at $(x_{i},y_{i})$. Since vorticity is the curl of the velocity, for the system of point vortices we have
@@
where we use complex notations $z=x+{\rm i}y$. In (3), the full velocity field $\mathbf{v}=(u,w)$ is the sum of the irrotational (i.e., zero-curl) background flow $\mathbf{V}=(U,W)$ and the flow created by superposition of $N$ point vortices. From (2, 3), $\mathbf{V}$ is a zero-divergence field. Since it is irrotational, it is also a gradient of a potential, and consequently, the potential is a harmonic function. In Sections 1-10, we will consider unrestricted flows in the whole plane with velocities bounded at infinity. The only harmonic functions satisfying above conditions are linear functions, so the background flow has constant velocity (uniform background flow). Using (3), we find the average velocity of the flow in a small disc centered at $z=z_{i}$, thus, obtaining the following equations of motion of vortices in the uniform background flow
@@
where $k$ is a complex constant, and the overbar denotes the complex conjugation.
The classification of the stationary patterns of vortices is then reduced to the study of solution of the system of $N$ algebraic equations
@@
This system also has a two-dimensional electrostatic interpretation: (5) is nothing but an equilibrium condition for $N$ point electric charges ${\cal Q}_{1},\dots,{\cal Q}_{N}$ interacting pairwise through a two-dimensional Coulomb (logarithmic) potential, and placed in the homogeneous electric field $\bar{k}$. The equilibrium configuration is a stationary point of the electrostatic energy
@@
The stationary patterns with $k\not=0$ are called translating equilibria (or translating configurations/patterns) because in the reference frame where the fluid rests at infinity, the whole pattern translates with constant velocity $(\text{Im}\,k,\text{Re}\,k)/(2\pi)$. Below, we mainly deal with equilibria without background flow, i.e., $k=0$ equilibrium configurations (“static configurations/patterns”):
@@
This system of equations is invariant under rigid motions and scaling of the plane. From the invariance of the stationary value of the electrostatic energy $E=2\sum_{i<j}{\cal Q}_{i}{\cal Q}_{j}\log|z_{i}-z_{j}|$ under scaling of the plane ($z_{i}\to cz_{i}$), it follows that the system has solutions only if
$\sum_{i<j}{\cal Q}_{i}{\cal Q}_{j}=0.$ (7)
O’Neil showed [39] that for almost all values of $Q_{i}$, satisfying the above condition, the number of distinct solutions to (6), modulo rotations, translations and scaling of the plane, is $(N-2)!$.
In the case when $k\not=0$ the translational invariance of the stationary value of energy $E=\sum_{i}{\cal Q}_{i}(kz_{i}+\bar{k}\bar{z}_{i})+2\sum_{i<j}{\cal Q}_{i}{\cal Q}_{j}\log|z_{i}-z_{j}|$ leads to the following (“neutrality”) condition
$\sum_{i}{\cal Q}_{i}=0.$ (8)
For almost all values of ${\cal Q}_{i}$ satisfying this condition, the number of distinct, modulo translations of the plane, solutions of (5) is $(N-1)!$ [39].
In other words, for generic ${\cal Q}_{i}$, satisfying (7) or (8), dimension of the solution set is zero. There are, however, “resonant” values of ${\cal Q}_{i}$, where this dimension is positive, i.e., solutions depend on non-trivial continuous complex parameters. This paper mainly deals with such resonant cases. Sections 12 and 13 are devoted to periodic (“vortex street”) versions of such systems.
We also briefly consider non-resonant cases. These are equilibrium patterns in non-uniform background flows, discussed in Section 11, or a class of periodic patterns, discussed at the beginning of Section 13 . Such patterns are generated by applying Darboux transformations to the “classical” configurations, which were previously obtained using the polynomial method [19, 32]. The “classical” configurations, first found by Stieltjes [45], are related to classical orthogonal polynomials, such as Hermite, Laguerre, and Jacobi polynomials (see, e.g., [5, 36]) [^1] Among the other configurations analyzed using the polynomial method, one can mention nested vortex polygons in rotating fluids [5, 6]. We do not discuss these here..
The polynomial method is a technique that uses polynomial solutions to differential equations to find and classify stationary configurations of point vortices. Specifically, vortex positions are linked to the roots of certain polynomials, enabling the study of equilibria and their connections to various polynomial systems.
Significant advancements in the field followed the works of Tkachenko [47] and Bartman [9]. In particular, Bartman observed connections between the equilibria of systems consisting of vortices of two species with a circulation ratio $-1$ and rational solutions of the KdV hierarchy.
In the next two sections, we review the application of the polynomial method to systems consisting of finite number of vortices of two species with the circulation ratio $-1$, or circulation ratio $-2$. For such systems, the polynomial method allows one to demonstrate, through a recursive procedure, the existence of infinite sequences of static configurations with increasing dimensions of the solution set [9, 12, 31]. These sequences exhaust all possible static configurations for such systems.
In the case of the circulation ratio $-1$, the recursive procedure can be reformulated in terms of transformations of the Schrodinger operator through factorization (Darboux transformations). Then, factorization method enables one to generate all possible translating configurations as well as to find the Wronskian representation of configurations. In more general setting, it enables the construction of families of vortex street patterns, and families of configurations in the non-uniform background flows. In cases where configurations are related to rational or soliton solutions of the KdV hierarchy ($\Lambda=1$ configurations of Section 3, and those of sections 10, 12) there exist alternatives to factorization method, e.g. $\tau$-function method.
In the case of the circulation ratio $-2$, static configurations can be generated through factorization of the third-order operators (considered in Sections 8 and 9) [35]. These configurations are related to rational solutions of the Sawada-Kotera hierarchy, and their Pfaffian representation can be obtained using the $\tau$-function method (Appendix 1).
We also review families of multi-parametric configurations of three species of vortices (discussed in Sections 6, 7, and 9). These families, initially found using polynomial method [^2] Actually, one of these families was first constructed through factorization, but in a different context [22]. [42], can be constructed through the factorization of either the Schrodinger or third-order operators. A new (vortex-street) generalization of one such family is obtained using the factorization method in Section 13.
@@ 2 Vortices, Polynomials and Rational Primitives
Let us first consider static equilibria of $N=l+m$ charges (or vortices) of two species with a circulation ratio $-\Lambda$. Without loss of generality we can set the values of the first $l$ charges to $-1$ and those of the remaining $m$ charges to $\Lambda$:
@@
Next, we introduce the following polynomials in $z$
@@
whose roots correspond to positions of charges of the first and second species respectively. Since we are dealing with two distinct species, $p$ and $q$ do not have common or multiple roots (For cases involving common/multiple roots, see e.g. [21]). The static equilibrium condition (6) for this system of two species is equivalent to the following bi-linear differential equation for the above polynomials
@@
where prime denotes differentiation with respect to $z$. One can check this equivalence by applying partial fraction decomposition to the ratio of the left-hand side of (10) and $pq$: Using the identity $\frac{1}{(z-z_{i})(z-z_{j})}=\frac{1}{z_{i}-z_{j}}\left(\frac{1}{z-z_{i}}-\frac{1}{z-z_{j}}\right)$ and rearranging terms, we obtain the following expression for the ratio [^3] For a more detailed derivation of poly-linear equations, see e.g. section 5.1 of [40].
$\sum_{i}\frac{2{\cal Q}_{i}\sum_{j\not=i}\frac{{\cal Q}_{j}}{z_{i}-z_{j}}}{z-z_{i}},\quad{\cal Q}_{i}\in\{-1,\Lambda\}.$
This expression vanishes identically in $z$ if and only if each residue $2{\cal Q}_{i}\sum_{j\not=i}\frac{{\cal Q}_{j}}{z_{i}-z_{j}}$ equals zero, i.e., if and only if $\sum_{j\not=i}\frac{{\cal Q}_{j}}{z_{i}-z_{j}}=0$. The latter is precisely the equilibrium conditions (6). Therefore (10) and (6) are equivalent when ${\cal Q}_{i}\in\{-1,\Lambda\}$.
The $\Lambda=1$ specification of (10)
$p^{\prime\prime}q-2p^{\prime}q^{\prime}+pq^{\prime\prime}=0,$ (11)
@@
is called the Tkachenko equation [47] in the literature on vortex patterns. Its generic solutions (i.e., those without common/multiple roots) describe equilibria of two species of vortices (charges) of equal magnitude but opposite sign. Bartman [9] identified the Tkachenko equation with the recurrence relation (17) for the Adler-Moser polynomials. As shown in the works of Burchnall and Chaundy [12] (see the next section), any generic polynomial solution of the Tkachenko equation is (modulo multiplication by arbitrary constants, and a shift of $z$) a pair of the Adler-Moser polynomials [1]:
$p(z)=P_{n\pm 1}(z),\quad q(z)=P_{n}(z).$ (12)
@@
In this way, we obtain a complete classification of static configuration of a system of vortices (charges) with strengths ±$\pm 1$. The degree of the $n$th Adler-Moser polynomial equals $n(n+1)/2$, which can be derived from general considerations without analyzing the Tkachenko equation. Indeed, static configurations exist only when the condition (7) holds. For charges ${\cal Q}_{i}=\pm 1$, this condition is satisfied only if $l$ and $m$ are consecutive triangular numbers. Therefore, equilibrium is only possible when the number of negative and positive charges (or vice versa) equals $l=n(n-1)/2$ and $m=n(n+1)/2$. The equilibrium positions of charges are the roots of consecutive Adler-Moser polynomials. These polynomials possess a non-trivial property: $P_{n}(z)$ is also a function of $n$ free parameters [^4] Including the shift of $z$. Without loss of generality, this trivial parameter can be set to zero and is excluded from the standard definition of Adler-Moser polynomials., so all possible equilibria of $N=l+m=n^{2}$ charges form an $n$-dimensional complex subspace of the $n^{2}$ dimensional complex space.
@@
Details on Adler-Moser polynomials will be provided in the following sections. Here, we note that these polynomials were first found by Burchnall and Chaundy [12], later re-discovered in the context of the Hadamard problem [30] by Lagnese and Stellmacher, and subsequently by Adler and Moser as polynomial $\tau$-functions corresponding to rational solutions of the KdV hierarchy of integrable PDEs [1]. Burchnall and Chaundy [12] constructed these polynomials as solutions of the Tkachenko equation (11) by considering it as a linear second order ODE for $p$, where coefficients of ODE are determined by some polynomial $P_{n}=q$. Then the requirement that its two linearly independent solutions $p=P_{n-1}$ and $p=P_{n+1}$ are both polynomial is equivalent to rationality in $z$ of two indefinite integrals (rational primitives)
@@
It turned out that, in addition to the Adler-Moser ($\Lambda=1$) case, there is another case with $\Lambda=2$, where an infinite chain of polynomial solutions to (10) exists and can be constructed with the help of rational primitives [9, 31].
@@ 3 $\Lambda=1$ and $\Lambda=2$ Chains: Main Sequences
where $C_{n}$ is a constant [^5] The values of such constants are typically chosen so that $p_{n}$ and $q_{n}$ are monic in $z$.. In order to build a recursive chain we require that both $p_{n-1}$ and $p_{n}$ be polynomial in $z$, so the primitive
@@
must be rational. Alternatively, by considering (10) as an ODE for $q$, we arrive at the requirement of rationality of
@@
It is clear that the pair of the rationality conditions above requires that both $2\Lambda$ and $2/\Lambda$ be integers, which is possible only if
@@
The case $\Lambda=1$ has already been mentioned and corresponds to the Adler-Moser polynomials. Since equation (10) is invariant under the involution $\Lambda\leftrightarrow 1/\Lambda$, $p\leftrightarrow q$, the cases $\Lambda=1/2$ and $\Lambda=2$ are equivalent. Therefore, without loss of generality, both can be represented by the $\Lambda=2$ case.
Now, let us take $p$ and $q$ that do not have multiple/common roots and factorize $p(z)$ as $p(z)=(z-z_{i})\tilde{p}(z)$, where $z_{i},i\leq l$ is a root of $p$ (see definition (9)). Then, residue of a simple pole at $z=z_{i}$ in (13) equals
$\frac{\partial}{\partial z}\left(\frac{q(z)^{2\Lambda}}{\tilde{p}(z)^{2}}\right)_{z=z_{i}}=\frac{q(z_{i})^{2\Lambda-1}}{\tilde{p}(z_{i})^{3}}(2\Lambda\tilde{p}(z_{i})q^{\prime}(z_{i})-2\tilde{p}^{\prime}(z_{i})q(z_{i})).$
Since $\tilde{p}(z_{i})=p^{\prime}(z_{i})$ and $\tilde{p}^{\prime}(z_{i})=p^{\prime\prime}(z_{i})/2$, the expression for the residue becomes
$-\frac{q(z_{i})^{2\Lambda-1}}{\tilde{p}(z_{i})^{3}}\left(p^{\prime\prime}(z_{i})q(z_{i})-2\Lambda p^{\prime}(z_{i})q^{\prime}(z_{i})\right).$
The last factor in the above expression vanishes, since $p$ and $q$ satisfy (10). Due to the absence of multiple roots, $\tilde{p}(z_{i})\not=0$, and the entire expression equals zero. Therefore, a simple pole is absent at $z=z_{i}$, and (13) is rational when (10) holds. By exchanging $p$ with $q$, $\Lambda$ with $1/\Lambda$, and repeating all the above arguments, we can demonstrate rationality of (14).
It is worth mentioning that the converse statement also holds, i.e. the bilinear equation (10) follows from the rationality of (13) and (14) (for details, see [31]).
Thus, once a pair of polynomials $p,q$ without multiple or common roots satisfying (10,15) is found, one can construct an infinite chain of polynomial solutions of (10) with the help of recursive procedure for $p_{n}$ and $q_{n}$ mentioned at the beginning of this section.
In more detail: Since $p=z^{l}+\dots$ and $q=z^{m}+\dots$, from the leading term of (10) we get the Diophantine equations [^6] These equations can be also obtained from (7). connecting $l$ and $m$
@@
These are quadratic equations, and to each $l$ there correspond two values of $m$ and vice versa (since to each $p$ there correspond two $q$’s and vice versa). In other words, sequence of solutions has the form
$\dots,(l_{i},m_{i}),(l_{i},m_{i+1}),(l_{i+1},m_{i+1}),\dots$
@@
Now, consider bilinear relation (10) with $p=p_{i}$ and $q=q_{i}$ as a linear second-order ODE with solution $q_{i}$. Its second linearly independent solution, $q=q_{i+1}$, is given by:
@@
Similarly, considering $p_{i-1}$ and $p_{i}$ as linearly independent solutions, we obtain:
@@
Due to freedom in choosing linearly independent solutions of (10), one can also write analogs of (19) and (20) for decreasing indices:
@@
Thus, one can generate $p_{i}$, and $q_{i}$ iteratively in either directions starting at some $i$.
Rewriting (19,20) or (21) in the differential form, we obtain the first-order differential recurrence relations:
@@
In the case $\Lambda=1$, the above recurrence relations become the first-order relations for the Adler-Moser polynomials $P_{n}$, with $q_{i}=P_{2i}$ and $p_{i}=P_{2i+1}$:
@@
Examples of several first Adler-Moser polynomials are :
$P_{0}=1,\quad P_{1}=z,\quad P_{2}=z^{3}+s_{1},\quad P_{3}=z^{6}+5s_{1}z^{3}+s_{2}z-5s_{1}^{2},\quad\dots,$ (24)
where $s_{i}$ are free parameters related bi-rationally to the KdV “times” [1, 3] (we set parameter related to shift of $z$ to zero). For more details on this relationship, see Appendix 1.
In the case $\Lambda=2$, we obtain the following recurrence relations: